# Young Diagram Decompositions for Almost Symmetric Numerical Semigroups

### v1

@@ 5. Decompositions of Almost Symmetric Numerical Semigroups
$R=S\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}S^{*}\text{ or }R=S\boxplus_{B}\{0,t+1,\rightarrow\}\boxplus_{C}S^{*}.$
@@
Since $R^{\prime}$ is symmetric, $\operatorname{g}(S)=n-k$. Then let $\operatorname{G}(S)=\{b_{1},\dots,b_{n-k}\}$, $f=F(S)=b_{n-k}=\frac{\operatorname{C}(R)-t-1}{2}$ and
@@
$S\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}S^{*}=$$\{0,r_{1},\dots,r_{k}^{\prime},\rightarrow\}\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}\{0,f-b_{n-k-1},\dots,f-b_{1},r_{k}^{\prime},\rightarrow\}$
@@
$\hskip 125.19194pt\dots,r_{k}^{\prime}+t-1+f-b_{1},r_{k}^{\prime}+t-1+r_{k}^{\prime},\rightarrow\}.$
@@
$S\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}S^{*}=$$\{0,r_{1},\dots,r_{k},\operatorname{F}(R^{\prime})-b_{n-k-1}+t-1,\dots,\operatorname{F}(R^{\prime})-b_{1}+t-1,\operatorname{C}(R),\rightarrow\}$
@@
$R=S\boxplus_{B}\{0,4,\rightarrow\}\boxplus_{C}S^{*},$
and the Young diagram of $Y_{R}=Y_{S}\boxplus_{C}Y_{\{0,4\rightarrow\}}\boxplus_{B}Y_{S^{*}}$ is depicted below
[…]
$\underbrace{\hskip 24.00002pt}_{\displaystyle Y_{S}}$ $\underbrace{\hskip 8.00003pt}_{\displaystyle Y_{\{0,4,\rightarrow\}}}$ $\underbrace{\hskip 55.99998pt}_{\displaystyle Y_{S^{*}}}$ $\underbrace{\hskip 80.0pt}_{\displaystyle Y_{R}=Y_{S}\boxplus_{C}Y_{\{0,4,\rightarrow\}}\boxplus_{B}Y_{S^{*}}}$
@@
$R=S\boxplus_{B}\{0,5,\rightarrow\}\boxplus_{C}S^{\star},$
and the Young diagram of $Y_{R}=Y_{S}\boxplus_{B}Y_{\{0,5\rightarrow\}}\boxplus_{C}Y_{S^{*}}$ is depicted below
[…]
$\underbrace{\hskip 15.99998pt}_{\displaystyle Y_{S}}$ $\underbrace{\hskip 8.00003pt}_{\displaystyle Y_{\{0,4,\rightarrow\}}}$ $\underbrace{\hskip 55.99998pt}_{\displaystyle Y_{S^{*}}}$ $\underbrace{\hskip 80.0pt}_{\displaystyle Y_{R}=Y_{S}\boxplus_{B}Y_{\{0,5,\rightarrow\}}\boxplus_{B}Y_{S^{*}}}$
@@
[12, cf. Theorem 4.10] Let $R=\{0,r_{1},\dots,r_{n},\rightarrow\}$ be a pseudo-symmetric numerical semigroup and $R\neq\{0,3,\rightarrow\}$. Then there is a unique numerical semigroup $S$ such that
@@
    $R\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}R^{*}=R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*}$ is a numerical semigroup if and only if $r_{n},r_{n}+1,\dots,r_{n}+t-2$ are minimal generators for $R$ where $t<r_{n-1}+r_{1}-r_{n}+2$, and $2r_{n}-r_{i}-r_{j}+t-2\neq r_{k}$ $\forall i,j,k\in[0,n-1]\cap\mathbb{Z}$.
@@
    $R\boxplus_{B}\{0,t+1,\rightarrow\}\boxplus_{C}R^{*}$ is a numerical semigroup if and only if $r_{n},r_{n}+1,\dots,r_{n}+t-1$ are minimal generators for $R$ where $t<r_{n-1}+r_{1}-r_{n}+1$, as well as $r_{n}-r_{i}+t-1\neq r_{k}$ and $2r_{n}-r_{i}-r_{j}+t-2\neq r_{k}$ $\forall i,j,k\in[0,n-1]\cap\mathbb{Z}$.
@@
    $R\boxplus_{C}\{0,t-1,\rightarrow\}\boxplus_{C}R^{*}$ is a numerical semigroup if and only if $r_{n},r_{n}+1,\dots,r_{n}+t-3$ are minimal generators for $R$ where $t<r_{n-1}+r_{1}-r_{n}+3$, and $2r_{n}-r_{i}-r_{j}+t-4\neq r_{k}$ $\forall i,j,k\in[0,n-1]\cap\mathbb{Z}$.
@@
By definition, $R\boxplus_{C}\{0,t,\rightarrow\}=\{0,r_{1},\dots,r_{n-1},r_{n}+t-1,\rightarrow\}$ and $\operatorname{G}(R\boxplus_{C}\{0,t,\rightarrow\})=\{a_{1},\dots,a_{g},r_{n},\dots,r_{n}+t-2\}$. Remember that $Y_{R}$ has $g$ rows and $n$ columns. Thus, the associated Young diagram $Y_{R\boxplus_{C}\{0,t,\rightarrow\}}$ of $R\boxplus_{C}\{0,t,\rightarrow\}$ is obtained by adding $t-1$ rows of length $n$ on top of $Y_{R}$. This corresponds to the $(t-1)\times n$ matrix shaped and orange part of Figure 5, and the hook lengths of boxes in this part are labeled with $u_{ij}$ where $i\in\{1,\dots,t-1\}$ and $j\in\{1,\dots,n\}$. Then by Proposition 3.2, we see that $u_{ij}=r_{n}-r_{j-1}+t-i-1$ for every $i$ and $j$.
@@
Therefore, the associated Young diagram $Y_{R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*}}$ of $R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*}$ is obtained by adding $n$ rows of length $n$ on top of $Y_{R\boxplus_{C}\{0,t,\rightarrow\}}$ and then glueing $Y_{R^{*}}$ to the right of it as shown in Figure 5. The hook lengths of boxes in the top left $n\times n$ matrix shaped and white part of Figure 5 are labeled with $\mu_{ij}$. Then by Proposition 3.2 again, we see that $\mu_{ij}=2r_{n}-r_{i-1}-r_{j-1}+t-2$ for every $i,j\in[1,n]\cap\mathbb{Z}$.
[…]
@@
This means that $u_{i1}=r_{n}+t-i-1$ can not be written as a positive factorization of $\{0,r_{1},\dots,r_{n-1}\}$, and so $u_{i1}$ is a minimal generator of $R$ for each $i\in[1,t-1]\cap\mathbb{Z}$. On the other hand, $u_{11}=r_{n}+t-2<r_{n-1}+r_{1}$ implies that $t<r_{n-1}+r_{1}-r_{n}+2$. Otherwise, one of $u_{i1}$ must be an element of $R\boxplus_{C}\{0,t,\rightarrow\}$, which is clearly a contradiction.
Further $\mu_{ij}\in\operatorname{G}(R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*})$ if and only if $2r_{n}-r_{i}-r_{j}+t-2\neq r_{k}$, $12r_{n}-r_{i}-r_{j}+t-2\neq r_{n}+t-$ 1 and $22r_{n}-r_{i}-r_{j}+t-2\neq 2r_{n}-a_{l}+t-$ 2 for every $i,j,k\in[0,n-1]\cap\mathbb{Z}$ and $l\in[1,g-1]\cap\mathbb{Z}$. However, since $R$ is a numerical semigroup, 2r_n-r_i-r_j+t-2≠r_n+t-1 ⇔r_n-1≠r_i+r_j and 2r_n-r_i-r_j+t-2≠2r_n-a_l+t-2 ⇔r_i+r_j≠a_l always hold for every $i,j,k\in[0,n-1]\cap\mathbb{Z}$ and $l\in[1,g-1]\cap\mathbb{Z}$. Hence, μ_ij∈G(R⊞_C{0,t,→}⊞_ER^*) ⇔2r_n-r_i-r_j+t-2≠r_k for every $i,j,k\in[0,n-1]\cap\mathbb{Z}$. This proves (1), and others follow in a similar way.

### v2

@@ 5. Decompositions of Almost Symmetric Numerical Semigroups
[…]
@@
Since $R^{\prime}$ is symmetric, $\operatorname{g}(S)=n-k$. Then let $\operatorname{G}(S)=\{b_{1},\dots,b_{n-k}\}$, $f=\operatorname{F}(S)=b_{n-k}=\frac{\operatorname{C}(R)-t-1}{2}$ and
@@
$S\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}S^{*}=$ $\{0,r_{1},\dots,r_{k}^{\prime},\rightarrow\}\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}$
$\hskip 68.28644pt\{0,f-b_{n-k-1},\dots,f-b_{1},r_{k}^{\prime},\rightarrow\}$
@@
$\hskip 68.28644pt\dots,r_{k}^{\prime}+t-1+f-b_{1},r_{k}^{\prime}+t-1+r_{k}^{\prime},\rightarrow\}.$
@@
$S\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}S^{*}=$ $\{0,r_{1},\dots,r_{k},\operatorname{F}(R^{\prime})-b_{n-k-1}+t-1,\dots$
$\hskip 68.28644pt\dots,\operatorname{F}(R^{\prime})-b_{1}+t-1,\operatorname{C}(R),\rightarrow\}$
@@
$R=S\boxplus_{C}\{0,4,\rightarrow\}\boxplus_{B}S^{*},$
and the Young diagram of $Y_{R}=Y_{S}\boxplus_{C}Y_{\{0,4,\rightarrow\}}\boxplus_{B}Y_{S^{*}}$ is depicted below
$\boxplus_{C}$ $\boxplus_{B}$ ==
$\underbrace{\hskip 35.3338pt}_{\displaystyle Y_{S}}$ $\underbrace{\hskip 11.77797pt}_{\displaystyle Y_{\{0,4,\rightarrow\}}}$ $\underbrace{\hskip 82.44548pt}_{\displaystyle Y_{S^{*}}}$ $\underbrace{\hskip 117.7793pt}_{\displaystyle Y_{R}=Y_{S}\boxplus_{C}Y_{\{0,4,\rightarrow\}}\boxplus_{B}Y_{S^{*}}}$
@@
$R=S\boxplus_{B}\{0,5,\rightarrow\}\boxplus_{C}S^{*},$
and the Young diagram of $Y_{R}=Y_{S}\boxplus_{B}Y_{\{0,5,\rightarrow\}}\boxplus_{C}Y_{S^{*}}$ is depicted below
$\boxplus_{B}$ $\boxplus_{C}$ ==
$\underbrace{\hskip 23.55583pt}_{\displaystyle Y_{S}}$ $\underbrace{\hskip 11.77797pt}_{\displaystyle Y_{\{0,5,\rightarrow\}}}$ $\underbrace{\hskip 82.44548pt}_{\displaystyle Y_{S^{*}}}$ $\underbrace{\hskip 117.7793pt}_{\displaystyle Y_{R}=Y_{S}\boxplus_{B}Y_{\{0,5,\rightarrow\}}\boxplus_{C}Y_{S^{*}}}$
@@
[13, cf. Theorem 4.10] Let $R=\{0,r_{1},\dots,r_{n},\rightarrow\}$ be a pseudo-symmetric numerical semigroup and $R\neq\{0,3,\rightarrow\}$. Then there is a unique numerical semigroup $S$ such that
@@
    $R\boxplus_{C}\{0,t+1,\rightarrow\}\boxplus_{B}R^{*}=R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*}$ is a numerical semigroup if and only if $r_{n},r_{n}+1,\dots,r_{n}+t-2$ are minimal generators for $R$ where $t<r_{n-1}+r_{1}-r_{n}+2$, and $2r_{n}-r_{i}-r_{j}+t-2\neq r_{k}$ $\forall i,j,k\in[0,n-1]\cap\mathbb{N}_{0}$.
@@
    $R\boxplus_{B}\{0,t+1,\rightarrow\}\boxplus_{C}R^{*}$ is a numerical semigroup if and only if $r_{n},r_{n}+1,\dots,r_{n}+t-1$ are minimal generators for $R$ where $t<r_{n-1}+r_{1}-r_{n}+1$, as well as $r_{n}-r_{i}+t-1\neq r_{k}$ and $2r_{n}-r_{i}-r_{j}+t-2\neq r_{k}$ $\forall i,j,k\in[0,n-1]\cap\mathbb{N}_{0}$.
@@
    $R\boxplus_{C}\{0,t-1,\rightarrow\}\boxplus_{C}R^{*}$ is a numerical semigroup if and only if $r_{n},r_{n}+1,\dots,r_{n}+t-3$ are minimal generators for $R$ where $t<r_{n-1}+r_{1}-r_{n}+3$, and $2r_{n}-r_{i}-r_{j}+t-4\neq r_{k}$ $\forall i,j,k\in[0,n-1]\cap\mathbb{N}_{0}$.
@@
By definition, $R\boxplus_{C}\{0,t,\rightarrow\}=\{0,r_{1},\dots,r_{n-1},r_{n}+t-1,\rightarrow\}$ and $\operatorname{G}(R\boxplus_{C}\{0,t,\rightarrow\})=\{a_{1},\dots,a_{g},r_{n},\dots,r_{n}+t-2\}$. Remember that $Y_{R}$ has $g$ rows and $n$ columns. Thus, the associated Young diagram $Y_{R\boxplus_{C}\{0,t,\rightarrow\}}$ of $R\boxplus_{C}\{0,t,\rightarrow\}$ is obtained by adding t1t-1 rows of length $n$ on top of $Y_{R}$. This corresponds to the $(t-1)\times n$ matrix shaped and orange part of Figure 1, and the hook lengths of boxes in this part are labeled with $u_{ij}$ where $i\in\{1,\dots,t-1\}$ and $j\in\{1,\dots,n\}$. Then by Proposition 3.2, we see that $u_{ij}=r_{n}-r_{j-1}+t-i-1$ for every $i$ and $j$.
@@
Therefore, the associated Young diagram $Y_{R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*}}$ of $R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*}$ is obtained by adding $n$ rows of length $n$ on top of $Y_{R\boxplus_{C}\{0,t,\rightarrow\}}$ and then glueing $Y_{R^{*}}$ to the right of it as shown in Figure 1. The hook lengths of boxes in the top left $n\times n$ matrix shaped and white part of Figure 1 are labeled with $\mu_{ij}$. Then by Proposition 3.2 again, we see that $\mu_{ij}=2r_{n}-r_{i-1}-r_{j-1}+t-2$ for every $i,j\in[1,n]\cap\mathbb{N}$.
$\mu_{11}$ $\mu_{1n}$ $\vdots$ $\vdots$ $\mu_{i1}$ $\mu_{in}$ $\vdots$ $\vdots$ $\mu_{n1}$ $\mu_{nn}$ $a_{g}$ $\dots$ $\vdots$ $\dots$ $a_{g}-r_{i-1}$ $\dots$ $\vdots$ $\dots$ $a_{g}-r_{n-1}$ $\dots$
$u_{11}$ $u_{1n}$ $\vdots$ $\vdots$ $u_{v1}$ $u_{vn}$
$a_{g}$ $a_{g}-r_{n-1}$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
@@
This means that $u_{i1}=r_{n}+t-i-1$ can not be written as a positive factorization of $\{0,r_{1},\dots,r_{n-1}\}$, and so $u_{i1}$ is a minimal generator of $R$ for each $i\in[1,t-1]\cap\mathbb{N}$. On the other hand, $u_{11}=r_{n}+t-2<r_{n-1}+r_{1}$ implies that $t<r_{n-1}+r_{1}-r_{n}+2$. Otherwise, one of $u_{i1}$ must be an element of $R\boxplus_{C}\{0,t,\rightarrow\}$, which is clearly a contradiction.
Further $\mu_{ij}\in\operatorname{G}(R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*})$ if and only if $2r_{n}-r_{i}-r_{j}+t-2\neq r_{k}$, $2r_{n}-r_{i}-r_{j}+t-2\neq r_{n}+t-1$ and $2r_{n}-r_{i}-r_{j}+t-2\neq 2r_{n}-a_{l}+t-2$ for every $i,j,k\in[0,n-1]\cap\mathbb{N}_{0}$ and $l\in[1,g-1]\cap\mathbb{N}$. However, since $R$ is a numerical semigroup,
$2r_{n}-r_{i}-r_{j}+t-2\neq r_{n}+t-1\iff r_{n}-1\neq r_{i}+r_{j}$
and
$2r_{n}-r_{i}-r_{j}+t-2\neq 2r_{n}-a_{l}+t-2\iff r_{i}+r_{j}\neq a_{l}$
always hold for every $i,j,k\in[0,n-1]\cap\mathbb{N}_{0}$ and $l\in[1,g-1]\cap\mathbb{N}$. Hence,
$\mu_{ij}\in\operatorname{G}(R\boxplus_{C}\{0,t,\rightarrow\}\boxplus_{E}R^{*})\iff 2r_{n}-r_{i}-r_{j}+t-2\neq r_{k}$
for every $i,j,k\in[0,n-1]\cap\mathbb{N}_{0}$. This proves (1), and others follow in a similar way.